# AMSbUmsbmn

### v1

@@ Abstract.
We conjecturally generalize the Conjecture by Briançon and Iarrobino on the most singular points of the Hilbert scheme of a tetrahedral number of points in ${\mathbb{A}}^{3}$ by introducing the notion of locally-maximal singularity. The formulation of this conjecture is based on the shape of the ideals via a picturesque pattern. It also generalizes the conjectural necessary condition for the most singular points, suggested by the second-named author in [24], restricted to a tetrahedral number of points in 3D. These conjectures and those in the prequel work aim to prove the original conjecture.
@@ 0. Introduction
Hilbert schemes of points are natural and crucial spaces in algebraic geometry. However, they are known to be highly singular. There has been tremendous growth in research to understand the singularities and behaviour of these spaces. For example, for Murphy’s law for the Hilbert scheme, see [27], and for a proof for the Hilbert schemes of points, see [13]; for the Hilbert scheme of points on threefolds, see [21, 15, 11, 16]; for the reducibility, see [12, 4, 9, 5]; for rational singularities, see [22]; for irrational components of the Hilbert scheme of points, see [6]; for the parity conjecture, see [19, 23, 7]; for other works related to enumerative geometry, see [1, 2, 14, 17, 18, 20, 25].
Because of the extremely singular nature of the Hilbert schemes, measuring the singularity is of great importance. Let $N\geq 3$. The dimension of the tangent space to $\mathrm{Hilb}^{l}({\mathbb{A}}^{N})$ at each point measures how singular the point is. If $l$ is a tetrahedral number (${N-1+k\choose N}$, Briançon and Iarrobino predicted the most singular points of the Hilbert scheme:
Conjecture A.
[3, Briançon-Iarrobino, 1978] The most singular point of $\operatorname{Hilb}^{{N-1+k\choose N}}({\mathbb{A}}^{N})$ corresponds to the ideal $\mathfrak{m}^{k}$, where $\mathfrak{m}$ is the maximal ideal in ${\mathbb{C}}[x_{1},x_{2},\dots,x_{N}]$.
This conjecture is still open. To pave the way towards proving this long-standing conjecture, investigating the relationship between the shape of the ideals corresponding to the points of the Hilbert scheme and their level of singularity was suggested in [24]. In this article, we further generalize the necessary condition in [24] using a novel and stunning pattern in the shape of the locally-maximal singularities (see Definition 1.1).
Let $A={\mathbb{C}}[x_{1},x_{2},\ldots,x_{N}]$. The statement of [24, Lemma 1.7] implies that if we prove the following necessary condition for being the most singular point of the Hilbert scheme ([24, Conjecture B]), then, in particular, we can resolve Conjecture A.
Conjecture B.
[24, Conjecture B] For $N\geq 3$, let $I$ be a 0-dimensional Borel-fixed ideal of colength $l$ in $A$ given by
@@
where $m_{1}\leq m_{2}\leq\ldots\leq m_{N}$.
If (${N-1+k\choose N}\leq l<{N+k\choose N}$, and $I$ correspond to the most singular point of $\mathrm{Hilb}^{l}({\mathbb{A}}^{N})$, then $m_{1}=k$.
We conjecturally extend the Briançon-Iarrobino conjecture on the maximal dimension of the tangent space to the Hilbert scheme of points in 3D. We propose a weaker version of this conjecture, which will still be enough to conclude Conjecture A.
@@
Apart from the notations above, for fixed (co)length $l$ and the smallest pure exponent $m_{1}$ (which, by the above convention, is the exponent of the first variable $x_{1}$), we denote by $T_{\max,m_{1}}(l)$, the maximal dimension of the tangent space to $\mathrm{Hilb}^{l}({\mathbb{A}}^{3})$, among points corresponding to the Bore-fixed ideals ideals of (co)length $l$ and the exponent of the first variable $m_{1}$. Also, we have
$T_{\max}(l)=\max\{T_{\max,1}(l),T_{\max,2}(l),\ldots,T_{\max,k}(l)\}.$
Terminology
When we say the dimension of the tangent space of an ideal $I$, we mean the dimension of the tangent space of the Hilbert scheme $\operatorname{Hilb}^{\operatorname{colength}(I)}({\mathbb{A}}^{2})$ at the zero locus of $I$.
Acknowledgements
This work is an outcome of the Cambridge Summer Research In Mathematics (SRIM) program 2024, where AA and ZZ were supervised by FR. The authors were supported by the ERC Advanced Grant MSAG. AA was also partially supported by St. John’s College (Cambridge) funding, and FR was also supported by a UKRI grant EP/X032779/1. The authors thank Mark Gross for supporting the project, and the SRIM for making this work possible. The authors thank Tony Iarrobino for useful comments. We also thank Oscar Selby for his help with the code in Appendix B. We acknowledge the use of GeoGebra [10] and Macaulay2 [8].
1. Conjectures
From now on, we restrict our attention to the case of $N=3$, and for simplicity, we denote the variables by $x,y,z$. For any Borel-fixed ideal $I=(x^{m_{1}},y^{m_{2}},z^{m_{3}},\text{mixed monomial generators})$, we assume that $m_{1}\leq m_{2}\leq m_{3}$. We recall that Borel-fixed ideals are among the candidates for obtaining the maximal dimension of the tangent space. First, we introduce a definition.
Fix $l={2+k\choose 3}$ and $m_{1}$. We say that a Borel-fixed ideal $I$ of colength $l$ has a locally-maximal singularity or $m_{1}$-maximal singularity if it has the maximal dimension tangent space among all the Borel-fixed ideals of colength $l={2+k\choose 3}$ and the exponent of $x$ being $m_{1}$. As mentioned in Notations, we denote such a maximal dimension by $T_{\max,m_{1}}(l)$.
To prove Conjecture B for tetrahedral numbers and $N=3$, we propose a weaker version:
Conjecture C.
Fix $l={2+k\choose 3}$. Then, the $T_{\mathrm{max},m_{1}}(l)-m_{1}$ diagram is increasing (Figure 1(a)). In particular, Conjecture B holds for tetrahedral numbers (hence, Conjecture A holds).
More precisely, we propose the following conjectural statement.
Conjecture D.
Fix $l={2+k\choose 3}$ and $2\leq m_{1}\leq k$. The following Borel-fixed ideal with $1+m_{1}(k-m_{1}+2)+{m_{1}\choose 2}$ (minimal) number of generators has a locally-maximal singularity.
$I=$ $(x^{m_{1}},y^{k},z^{m_{3}},$
all degree k mixed generators xαyβzγ with 0αm11 and 0βk1 , $\text{all degree $k$ mixed generators $x^{\alpha}y^{\beta}z^{\gamma}$ with $0\leq\alpha\leq m_{1}-1$ and $0\leq\beta\leq k-1$ , $0\leq\gamma\leq m_{1}$},$
all degree k+1 mixed generators xαyβzγ with 0αm11 and $\text{all degree $k+1$ mixed generators $x^{\alpha}y^{\beta}z^{\gamma}$ with $0\leq\alpha\leq m_{1}-1$ and $\gamma=m_{1}+2<m_{3}$},$
all degree k+2 mixed generators xαyβzγ with 0αm11 and $\text{all degree $k+2$ mixed generators $x^{\alpha}y^{\beta}z^{\gamma}$ with $0\leq\alpha\leq m_{1}-1$ and $\gamma=m_{1}+4<m_{3}$},$
$\ldots$,… ,
all degree k+h mixed generators xαyβzγ with 0αm11 and $\text{all degree $k+h$ mixed generators $x^{\alpha}y^{\beta}z^{\gamma}$ with $0\leq\alpha\leq m_{1}-1$ and $\gamma=m_{1}+2h<m_{3}$},$
$\ldots$,… ,
$xz^{m_{1}+2(k-m_{1}-1)},yz^{m_{1}+2(k-m_{1}-1)}),$
where
(1.1) $m_{3}=l-\Big{(}m_{1}\big{(}{m_{1}+1\choose 2}+\sum_{i=1}^{k-2m_{1}+1}(m_{1}+2i)\big{)}+\sum_{j=1}^{m_{1}-2}(m_{1}-j)(m_{1}+2(k-2m_{1}+j+1))\Big{)},$
if $k\geq 2m_{1}-1$, and
(1.2) $m_{3}=l-\Big{(}m_{1}{m_{1}+1\choose 2}-{f\choose 2}m_{1}+{f\choose 3}+\sum_{i=0}^{k-m_{1}-2}(k-m_{1}-i)(m_{1}+2+2i)\Big{)},$
if $k\geq m_{1}\geq k-\left\lfloor\frac{k}{2}\right\rfloor$, where we define $f:=2m_{1}-k$, for $0\leq f\leq 2m_{1}$.
In the former case, the dimension of the tangent space to $\mathrm{Hilb}^{l}(\mathbb{A}^{3})$ at $I$ is given by
$T_{\max,m_{1}}(l)=3l+2{m_{1}\choose 2}\Big{(}2(k-m_{1}+1)^{2}+{m_{1}\choose 2}-1\Big{)},$
and in the latter case, if we define $g:=k-m_{1}$, for $0\leq g\leq\left\lfloor\frac{k}{2}\right\rfloor$, the dimension of the tangent space to $\mathrm{Hilb}^{l}(\mathbb{A}^{3})$ at $I$ is given by
$T_{\max,m_{1}}(l)={k+1\choose 2}{k+2\choose 2}-{g+1\choose 2}(k-g)(k+g+3).$
Remark 1.2.
Note that if $k=2m_{1}$ or $k=2m_{1}-1$, the two cases in Conjecture D coincide.
Remark 1.3.
Let $l={2+k\choose 3}$. For simplicity and to avoid separating into two cases, we assume $k$ is large enough, and we have the following graph that illustrates both of the conjectures above.
(a) *
Figure 1. Illustration of Conjecture C and Conjecture D for sufficiently large $k$. (note that the scales of the vertical and horizontal axes are chosen differently to be able to depict it here). The graph could be locally concave or convex; our conjectures only claim that the graph is increasing, which suffices to conclude Conjecture A in 3D and Conjecture B in 3D and for tetrahedral numbers.
[.99 ]Refer to caption
Remark 1.4.
We expect similar conjectures for arbitrary $N$. However, we focus only on $N=3$ throughout the paper to facilitate visualization.
2. Examples
@@
In this section, we verify the conjectures in some examples. In Appendix A, we provide the codes we used to compute examples in this section.
@@ 2.1. Table
As evidence, in the following table, we confirm Conjecture C and Conjecture D for $k\leq 5$. We have also confirmed the conjectures for a good portion of the Borel-fixed ideal for $k=6$. For the complete check for higher $k$, one may need to use Macaulay2 on high-performance computing systems. Already, for the cases that we have checked, the conjectural pattern is clear.
We analyze one of the examples in the table more precisely. For $k=5$, $l={k+2\choose 3}=35$, and $m_{1}=3$, the ideal
$(x^{3},y^{5},z^{7},y^{2}z^{3},xyz^{3},x^{2}z^{3},y^{3}z^{2},xy^{2}z^{2},x^{2}yz^{2},y^{4}z,xy^{3}z,x^{2}y^{2}z,xy^{4},x^{2}y^{3},yz^{5},xz^{5})$
is the unique Borel-fixed ideal of tangent space dimension $225$ among those with such length and $m_{1}$ by computation.
For this example, let us verify Conjecture D. Here $m_{2}=k=5$ as claimed, $m_{3}=7$ as yielded by (1.2). Note that in this case, $f=1$, and ($f\choose 2$ and ($f\choose 3$ are defined to be zero, as suggested by the combinatorial interpretation. There are $16$ generators as predicted and all mixed generators except the last two, $xz^{m_{1}+2(k-m_{1}-1)}$ and $yz^{m_{1}+2(k-m_{1}-1)}$, are of degree $5$. We check that the $x$-, $y$-, and $z$-exponents are bounded by $m_{1}-1=2$, $k-1=4$, and $m_{1}=3$, respectively. Moreover, the tangent space dimension is also predicted by the formula from the latter case.
length ($l$) k $m_{1}$ Borel-fixed ideal with $m_{1}$-maximal singularity $T_{\max,m_{1}}(l)$
4 2 2 $(x,y,z)^{2}$ 18
10 3 2 $(x^{2},y^{3},z^{4},xz^{2},yz^{2},y^{2}z,xy^{2},xyz)$ 46
3 $(x,y,z)^{3}$ 60
20 4 2 $(x^{2},y^{4},z^{6},y^{2}z^{2},xyz^{2},y^{3}z,xy^{2}z,xy^{3},yz^{4},xz^{4})$ 96
3 $(x^{3},y^{4},z^{5},yz^{3},xz^{3},y^{2}z^{2},xyz^{2},x^{2}z^{2},y^{3}z,xy^{2}z,x^{2}yz,xy^{3},$ 126
$x^{2}y^{2})$
4 $(x,y,z)^{4}$ 150
35 5 2 $(x^{2},y^{5},z^{9},y^{3}z^{2},xy^{2}z^{2},y^{4}z,xy^{3}z,xy^{4},y^{2}z^{4},xyz^{4},yz^{6},xz^{6})$ 169
3 $(x^{3},y^{5},z^{7},y^{2}z^{3},xyz^{3},x^{2}z^{3},y^{3}z^{2},xy^{2}z^{2},x^{2}yz^{2},y^{4}z,xy^{3}z,$ 225
$x^{2}y^{2}z,xy^{4},x^{2}y^{3},yz^{5},xz^{5})$
4 $(x^{4},y^{5},z^{6},xz^{4},yz^{4},y^{2}z^{3},xyz^{3},x^{2}z^{3},y^{3}z^{2},xy^{2}z^{2},x^{2}yz^{2},$ 279
$x^{3}z^{2},y^{4}z,xy^{3}z,x^{2}y^{2}z,x^{3}yz,xy^{4},x^{2}y^{3},x^{3}y^{2})$
5 $(x,y,z)^{5}$ 315
2.2. Shape of the locally-maximal singularities
Starting from the generators of the ideal $(x,y,z)^{k}$, which was predicted to be the most singular point of the Hilbert scheme of (${k+2\choose 3}$ points in ${\mathbb{A}}^{3}$, we predict the shape of Borel-fixed ideals with $m_{1}$-maximal singularities. We precisely predict the degree $k$ generators that remain on the triangle of generators of $(x,y,z)^{k}$ , and the higher or lower degree generators that do not stay on the triangle. Note that $y^{k}$ always remains on the triangle, and $z^{m_{3}}$ is always above the triangle. This is given in the statement of Conjecture D, but to clarify further, we give an explicit example and visualize the pattern with pictures.
Example 2.1.
Let $l=35$. The configuration of the generators of the Borel-fixed ideals of colength $35$ with maximal $T_{\max,m_{1}}(35)$ is depicted in Figure 2(a). The corresponding ideals are explicitly described in the table above. The generators of $(x,y,z)^{5}$ appear in all configurations for comparison, and the new generators appear in the colour of the surface of the generators. Also, we use a lighter colour for those generators of $(x,y,z)^{5}$ that are not the generators of the corresponding configuration.
(a) *
(b) *
(c) *
(d) *
Figure 2. Illustration of the configuration of the generators of the ideals with maximal dimension tangent space with fixed $2\leq m_{1}\leq 5=k$ and $l=35$ in Example 2.1. The pattern in the degeneration of the full triangle, as we decrease $m_{1}$ from $k$ to $2$, is clear in this picture. Note that $m_{2}$ is always fixed during degeneration, but $m_{3}$ increases by Formulas (1.1) and (1.2) as we decrease $m_{1}$. The corresponding ideals are (i) $(x,y,z)^{5}$, (ii) $(x^{4},y^{5},z^{6},xz^{4},yz^{4},y^{2}z^{3},xyz^{3},x^{2}z^{3},y^{3}z^{2},xy^{2}z^{2},x^{2}yz^{2},x^{3}z^{2},y^{4}z,xy^{3}z,\\ x^{2}y^{2}z,x^{3}yz,xy^{4},x^{2}y^{3},x^{3}y^{2})$, (iii) $(x^{3},y^{5},z^{7},y^{2}z^{3},xyz^{3},x^{2}z^{3},y^{3}z^{2},xy^{2}z^{2},\\ x^{2}yz^{2},y^{4}z,xy^{3}z,x^{2}y^{2}z,xy^{4},x^{2}y^{3},yz^{5},xz^{5})$ , and (iv) $(x^{2},y^{5},z^{9},y^{3}z^{2},xy^{2}z^{2},y^{4}z,xy^{3}z,xy^{4},y^{2}z^{4},xyz^{4},yz^{6},xz^{6})$.
[.38] Refer to caption                [.37] Refer to caption                     [.35] Refer to caption                [.36] Refer to caption
@@ Appendix A Macaulay2 codes
The codes we use are as follows. One of the most effective ways is to fix the colenght and also the number of (minimal) generators (from $22k+$ 2 to (${k+2\choose 2}$) and list all the Borel-fixed ideals and the dimension of the tangent space at the corresponding point of the Hilbert scheme. In the following codes, we fix $\operatorname{colength}=35$ and the number of generators, $18$. The third line lists all Borel-fixed ideals $I$ of colength $35$ with $18$ (minimal) generators along with the dimension of the tangent space to the Hilbert scheme at the point $[I]$.
@@ Appendix B A potential approach to prove the conjectures
Let $I$ be a Borel-fixed ideal of colength $l$ in ${\mathbb{C}}[x,y,z]$. Also, let $G$ be the (minimal) number of generators of $I$. By definition of the tangent space as $\operatorname{Hom}(I,R/I)$, to explicitly count the tangent vectors, one approach is to eliminate the vectors in the $Gl$-space of the potential maps from the generators of $I$ to $R/I$, which are not in the tangent space. We call such vectors zero vectors. Spotting patterns among the set of zero vectors for fixed $k$, $l={k+2\choose 3}$ and $m_{1}$ (the exponent of $x$, with the convention of being the smallest pure exponent, as in the first paragraph of Section 1) will lead to proving Conjecture C and Conjecture D.
@@
Let $I=(x,y,z^{2})^{2}$. Using the codes, we can list the zero vectors illustrated in Figure 3(a). Recall that this ideal has colenght $8$ and it has the maximal dimension tangent space, as considered in [26]. Using the zero vectors approach, we can calculate the dimension of the tangent space, $6\times 8-12=36$.
@@
Here, we provide some codes (in Python) that help list the zero vectors of any Borel-fixed ideal in three variables. Using the approach in [21], one can recognize the tangent vectors in 3D and then look at the complements as zero vectors. The following example calculates the dimension of the graded component of $T(I)$ of degree $\alpha\in\mathbb{Z}^{3}$ when $I=(x^{2},y^{3},z^{3},xy,xz,yz^{2},y^{2}z)$ and $\alpha=(0,2,-3)$.

### v2

@@ Abstract.
The Briançon-Iarrobino conjecture predicts the maximum singularity of the Hilbert scheme of a tetrahedral number of points. As for the maximal singularities of the Hilbert scheme of a non-tetrahedral number of points, the second named author gave some separate conjectural necessary and sufficient conditions. In this paper, we provide a conjectural sufficient condition for the necessary condition, and propose a monotonicity conjecture which predicts that for a fixed colength $l$, the maximal dimension of the tangent space over all the Borel-fixed ideals of colength $l$ is increasing with respect to the smallest pure exponent of the ideal.
@@ 0. Introduction
Hilbert schemes of points are natural and crucial spaces in algebraic geometry. However, they are known to be highly singular. There has been tremendous growth in research to understand the singularities and behavior of these spaces. For example, for Murphy’s law for the Hilbert scheme, see [Vakil06], and for a proof of a variant for the Hilbert schemes of points, see [Jelisiejew20]; for the Hilbert scheme of points on threefolds, see [Ramkumar-Sammartano, Jelisiejew-Ramkumar-Sammartano24, HuX, KatzS]; for the reducibility, see [Iarrobino72, Hilb8, Jardim, Hilb_11]; for rational singularities, see [Ramkumar-Sammartano24]; for irrational components of the Hilbert scheme of points, see [Farkas-Pandharipande-Sammartano24]; for the parity conjecture, see [PandharipandeSlides, Ramkumar-Sammartano1, GGGL23]; for other works related to enumerative geometry, see [BBS, BFHilb, JKSCounterBehrend, MNOP1, Nesterov25, PT, RicolfiSign].
Because of the extremely singular nature of the Hilbert schemes, measuring the singularity is of great importance. Let $N\geq 3$. The dimension of the tangent space to $\mathrm{Hilb}^{l}({\mdmathbb{A}}^{N})$ at each point measures how singular the point is. If $l$ is a tetrahedral number $N-1+k\choose N$}, Briançon and Iarrobino predicted the most singular points of the Hilbert scheme, which was recently resolved in three dimensions in [Mackenzie-Rezaee1]:
Conjecture A ([Bri-Iar]).
The most singular point of $\operatorname{Hilb}^{{N-1+k\choose N}}({\mdmathbb{A}}^{N})$ corresponds to the ideal $\mathfrak{m}^{k}$, where $\mathfrak{m}$ is the maximal ideal in ${\mdmathbb{C}}[x_{1},x_{2},\dots,x_{N}]$.
To generalize this to non-tetrahedral numbers, investigating the relationship between the shape of the ideals corresponding to the points of the Hilbert scheme and their level of singularity was suggested in [Rezaee-23-Conjectures]. In this article, we further generalize the necessary condition in [Rezaee-23-Conjectures] using a novel pattern in the shape of the locally-maximal singularities (see Definition 0.1).
First, we recall the necessary condition.
Conjecture B (Necessary condition for maximal singularity).
[Rezaee-23-Conjectures, Conjecture B] Let $A={\mdmathbb{C}}[x_{1},x_{2},\ldots,x_{N}]$. For $N\geq 3$, let $I$ be a 0-dimensional Borel-fixed ideal of colength $l$ in $A$ given by
@@
where $m_{1}\leq m_{2}\leq\ldots\leq m_{N}$. If ${N-1+k\choose N}\leq l<{N+k\choose N}$, and $I$ correspond to the most singular point of $\mathrm{Hilb}^{l}({\mdmathbb{A}}^{N})$, then $m_{1}=k$.
We propose a stronger version of this conjecture. Before stating the conjecture, we introduce some notations and definitions.
For any Borel-fixed ideal $I=(x_{1}^{m_{1}},x_{2}^{m_{2}},\ldots,x_{N}^{m_{N}},\text{mixed terms})$, we have $m_{1}\leq m_{2}\leq\ldots\leq m_{N}$. We recall that Borel-fixed ideals are among the candidates for obtaining the maximal dimension of the tangent space.
[…]
Now, we have the following main statements.
Conjecture C (Monotonicity conjecture).
For fixed $l$, the function $T_{\mathrm{max},m_{1}}(l)$ is increasing in $m_{1}$. In particular, Conjecture B holds.
See Figure 1(a).
(a) *
Figure 1. For fixed $l$, $T_{\mathrm{max},m_{1}}(l)$ is increasing in $m_{1}$.
[.98] Refer to caption
Acknowledgements
This work is an outcome of the Cambridge Summer Research In Mathematics (SRIM) program 2024, where AA and ZZ were supervised by FR. The authors were supported by the ERC Advanced Grant MSAG. AA was also partially supported by St. John’s College (Cambridge) funding, and FR was also supported by a UKRI grant EP/X032779/1. The authors thank Mark Gross for supporting the project, and the SRIM for making this work possible. The authors also thank Tony Iarrobino and Owen Mackenzie for helpful comments on the first version. We also thank Oscar Selby for his help with the code in Appendix B. We acknowledge the use of Macaulay2 [M2] and Python.
@@ Acknowledgements
We have the following maicn notations.
$T(I)$: The dimension of the tangent space to the Hilbert scheme at the ideal $I$,
which is defined by $\mathrm{hom}(I,R/{I})$.
$T_{\max,m_{1}}(l)$: The maximal dimension of the tangent space for fixed $l$ and $m_{1}$
(see Definition 0.1).
1. The monotonicity conjecture in examples
@@
In this brief section, we verify Conjecture C in some examples in the following table. In Appendix A, we provide the codes we used to compute examples in this section.
@@ 1.1. Table
As evidence, in the following table, we confirm Conjecture C for $N=3$ and $10\leq l\leq 35$. Note that for $N=3$ and $m_{1}=1$, we have $T_{\max,m_{1}}(l)=3l$ (so this case is not included in the table).
length ($l$) k $m_{1}$ $T_{\max,m_{1}}(l)$ length ($l$) k $m_{1}$ $T_{\max,m_{1}}(l)$ length ($l$) k $m_{1}$ $T_{\max,m_{1}}(l)$
10 3 2 46 21 4 2 99 4 190
3 60 3 129 29 4 2 141
11 3 2 49 4 153 3 183
3 63 22 4 2 104 4 207
12 3 2 54 3 132 30 4 2 144
3 66 4 156 3 186
13 3 2 61 23 4 2 109 4 210
3 69 3 141 31 4 2 149
14 3 2 64 4 159 3 207
3 78 24 4 2 114 4 217
15 3 2 69 3 148 32 4 2 156
3 81 4 168 3 210
16 3 2 78 25 4 2 123 4 240
3 88 3 151 33 4 2 159
17 3 2 81 4 177 3 213
3 99 26 4 2 126 4 243
18 3 2 86 3 158 34 4 2 166
3 102 4 180 3 216
19 3 2 89 27 4 2 129 4 276
3 123 3 161 35 5 2 169
20 4 2 96 4 187 3 225
3 126 28 4 2 134 4 279
4 150 3 176 5 315
@@ Appendix A Macaulay2 codes
The codes we use are as follows. One of the most effective ways is to fix the colenght and also the number of (minimal) generators (from $2k+2$ to $k+2\choose 2$}) and list all the Borel-fixed ideals and the dimension of the tangent space at the corresponding point of the Hilbert scheme. In the following codes, we fix $\operatorname{colength}=35$ and the number of generators, $18$. The third line lists all Borel-fixed ideals $I$ of colength $35$ with $18$ (minimal) generators along with the dimension of the tangent space to the Hilbert scheme at the point $[I]$.
@@ Appendix B A potential approach to prove the conjectures
Let $I$ be a Borel-fixed ideal of colength $l$ in ${\mdmathbb{C}}[x,y,z]$. Also, let $G$ be the (minimal) number of generators of $I$. By definition of the tangent space as $\operatorname{Hom}(I,R/I)$, to explicitly count the tangent vectors, one approach is to eliminate the vectors in the $Gl$-space of the potential maps from the generators of $I$ to $R/I$, which are not in the tangent space. We call such vectors zero vectors. Spotting patterns among the set of zero vectors for fixed $k$, $l={k+2\choose 3}$ and $m_{1}$ (the exponent of $x$, with the convention of being the smallest pure exponent, as in the first paragraph of Section 1) will lead to proving Conjecture C.
@@
Let $I=(x,y,z^{2})^{2}$. Using the codes, we can list the zero vectors illustrated in Figure 2(a). Recall that this ideal has colenght $8$ and it has the maximal dimension tangent space, as considered in [Sturmfels]. Using the zero vectors approach, we can calculate the dimension of the tangent space, $6\times 8-12=36$.
@@
[…]